the current in an electron beam in a cathode-ray tube is measured to be 80microA.The number of electrons that hit the screen in 8s is.

Grade:12

1 Answers

Khimraj
5 years ago
I = q/t = ne/t
so n = It/e
n = 80*10-6*8/1.6*10-19
n = 4*1015
Hope itclears.............................................................................................